OpenStudy (anonymous):

Pls help. Pls help. Solve the equation by using the quadratic formula. ((5)/(8))x - 8x + (5)/(2)=0 yesterdayReport Abuse

OpenStudy (anonymous):

20/59

OpenStudy (anonymous):

For the quadratic equation ax² + bx + c = 0 The quadratic formula is . . . . . . . . . _______ . . . .− b ±√b² − 4ac x = --------------------- . . . . . . . . 2a Now substituting the values . . . . . .___________ − 8 ±√(8)² − 4(1)(−5) --------------------------- . . . . . . . 2(1) . . . . . .______ − 8 ±√64 + 20 --------------------- . . . . . . 2 . . . . . .__ − 8 ±√84 ------------- . . . . 2 . . . . . . . __ − 8 ± 2√21 ---------------- . . . . 2 . . . . . . . . . .__ [ x = − 4 ±√21 ]

OpenStudy (anonymous):

\[\frac{5 x}{8}-8x+\frac{5}{2}\text{=}0\]Multiply each side by -8 to get rid of the denominators \[-8\left(\frac{5 x}{8}-8x+\frac{5}{2}\right)=-8*0 \]and then simplify.\[-20+59 x=0\]Put -20 on the right side and divide both sides by 59 to yield\[x=\frac{20}{59} \]penguinbaby's first post had the correct answer.